OpenStudy (anonymous):

What is the solution to the system of equations? x = 3y + 5 x - 4y = 3

OpenStudy (anonymous):

how about now

pooja195 (pooja195):

Ok we are given 2 equations x = 3y + 5 <---equation 1 x - 4y = 3 <---equation 2 We are given the value of x already so take equation 1 and plug it into equation 2 :) 3y+5-4y=3 solve for y

OpenStudy (okdutchman7):

You can substitute the value for x 4y+3=3y+5 y=2

pooja195 (pooja195):

@okdutchman7 I would appreciate it if you let the asker do the work. :3

OpenStudy (okdutchman7):

Plug y in and x= 4(2) + 3 x = 11

OpenStudy (anonymous):

thanks you guys

pooja195 (pooja195):

I hope you actually understood. And learned.

OpenStudy (anonymous):

so my answer is (11,2)

pooja195 (pooja195):

Yeah.

OpenStudy (anonymous):

lol thanks

pooja195 (pooja195):

yw

OpenStudy (anonymous):

i have a few more i'll tag you

pooja195 (pooja195):

Ok.